# Commutation of Smyth and Hoare Power Constructions in Well-filtered Dcpos

### v3

@@ 5 The double power well-filtered dcpo
During the study of monads, there has always been a problem worth thinking about: Given two monads with underlying functors $S$ and $T$ on a category $\mathbf{C}$, does $T\circ S$ (or $S\circ T$) still carry the structure of a monad? Until 1969, Beck introduced the concepts of distributive laws and showed that the existence of distributive laws provides a sufficient condition for such a composition to form a monad [2].
@@
([2]) Let $\mathcal{S}=(\mathcal{S},\eta^{\mathcal{S}},\mu^{\mathcal{S}})$ and $\mathcal{T}=(\mathcal{T},\eta^{\mathcal{T}},\mu^{\mathcal{T}})$ be two monads on the category $\mathbf{C}$. A distributive law of $\mathcal{S}$ over $\mathcal{T}$ is a natural transformation $l:\mathcal{TS}\rightarrow\mathcal{ST}$ such that the following diagrams commute (Figure 1).
@@
([1]) Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ and $\mathcal{T}^{\prime}=(\mathcal{T}^{\prime},\eta^{\prime},\mu^{\prime})$ be two monads. A morphism of monads is a natural transformation $\lambda:\mathcal{T}\rightarrow\mathcal{T}^{\prime}$ such that the following two diagrams commute (Figure 2).
@@
Particularly, the morphisms of monads are called triple maps by Beck in [2]. Meanwhile, Beck also proved the following theorem.
@@
([2]) Let $\mathcal{S}=(\mathcal{S},\eta^{\mathcal{S}},\mu^{\mathcal{S}})$, $\mathcal{T}=(\mathcal{T},\eta^{\mathcal{T}},\mu^{\mathcal{T}})$ be two monads on the category $\mathbf{C}$. Then the following statements are equivalent:
@@
      $(\mathcal{TS},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\eta_{\mathcal{S}}^{\mathcal{T}}\eta^{\mathcal{S}}},\mu)$ is a monad;
@@
  3. 3.
    Liftings of $\tilde{\mathcal{S}}$ of the monad $\mathcal{S}$ into $\mathbf{C}^{\mathcal{T}}$.
From the proposition presented in Section 2 of Beck [2], a more general result concerning the morphisms of monads was established by Fritz et al.
@@
([7]) Let $\mathcal{S}=(\mathcal{S},\eta_{\mathcal{S}},\mu_{\mathcal{S}})$ and $\mathcal{T}=(\mathcal{T},\eta_{\mathcal{T}},\mu_{\mathcal{T}})$ be two monads on a category $\mathbf{C}$ and $\lambda:\mathcal{S}\rightarrow\mathcal{T}$ a morphism of monads. Then every $\mathcal{T}$-algebra $(A,\alpha)$ can be equipped with an $\mathcal{S}$-algebra via $(A,\alpha)\mapsto(A,\alpha\circ\lambda)$. Moreover, a $\mathcal{T}$-algebra homomorphism $f:(A,\alpha)\rightarrow(B,\beta)$ in duces an $\mathcal{S}$-algebra homomorphism $f:(A,\alpha\circ\lambda)\rightarrow(B,\beta\circ\lambda)$ in a functorial way.
@@
Assume $L$ and $M$ are well-filtered dcpos and $f:L\rightarrow M$ is a Scott-continuous mapping. We need to show the following diagram commutes,
@@
First, suppose that there exists a $cl(f(A))\in\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})$ but $cl(f(A))\notin\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$, that is, we can find a $N_{0}\in cl\{{\uparrow}f(K):K\in\mathcal{A}\}$ satisfying $cl(f(A))\cap N_{0}=\emptyset$. From which we can see that $N_{0}\in\Box(M\setminus cl(f(A)))$. Since the well-filteredness of $M$ implies that $\Box(M\setminus cl(f(A)))\in\sigma(\mathcal{Q}(M))$, $\Box(M\setminus cl(f(A)))\cap\{{\uparrow}f(K):K\in\mathcal{A}\}\neq\emptyset$. So there is a $K_{0}\in\mathcal{A}$ such that ${\uparrow}f(K_{0})\subseteq M\setminus cl(f(A))$, that is to say, ${\uparrow}f(K_{0})\cap cl(f(A))=\emptyset$. Then $K_{0}\cap A=\emptyset$, which obviously contradicts the fact that $A$ meets every element in $\mathcal{A}$. Thus, $\{cl(f(A)):\forall K\in\mathcal{A},A\cap K\neq\emptyset\}\subseteq\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$; furthermore, $\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})\subseteq\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$.
Now we prove the reverse inclusion. Let $B\in\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$. Then for each $K\in\mathcal{A}$, $B\cap{\uparrow}f(K)\neq\emptyset$, combing with the downness of $B$, which further indicates that $B\cap f(K)\neq\emptyset$. We pick $x_{K}\in K$ such that $f(x_{K})\in B\cap f(K)$ and set $A_{0}=cl(\{x_{K}:K\in\mathcal{A}\})$. Then $cl(f(A_{0}))=cl(f(\{x_{K}:K\in\mathcal{A}\}))=cl(\{f(x_{K}):K\in\mathcal{A}\})\subseteq B$ by the closeness of $B$. Meanwhile, $A_{0}\cap K\neq\emptyset$ for each $K\in\mathcal{A}$. So $B\in\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})$, then, $\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})\subseteq\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})$ holds.
@@
First, suppose that there is a $K\in cl(\{{\uparrow}x:x\in A\})$ such that $A\cap K=\emptyset$, then $K\subseteq L\setminus A$, i.e., $K\in\Box(L\setminus A)$. By the well-filteredness of $L$, we know $\Box(L\setminus A)\in\sigma(\mathcal{Q}(L))$. So $\Box(L\setminus A)\cap\{{\uparrow}x:x\in A\}\neq\emptyset$, that is to say, we can find an $x\in A$ such that ${\uparrow}x\subseteq L\setminus A$. Thus $x\notin A$, clearly, which is a contradiction. So for any $K\in cl(\{{\uparrow}x:x\in A\}),A\cap K\neq\emptyset$, which implies that $A\in\phi_{L}\circ\mathcal{H}\theta_{L}(A)$; hence, $\theta_{\mathcal{H}(L)}(A)={\uparrow}\{A\}\subseteq\phi_{L}\circ\mathcal{H}\theta_{L}(A)$.
Now consider the reverse inclusion. For any $B\in\phi_{L}\circ\mathcal{H}\theta_{L}(A)$, we have $B\cap{\uparrow}x\neq\emptyset,\forall x\in A$. It means that each $x\in A$ belongs to $B$. Thus $A\subseteq B$, equivalently, $B\in{\uparrow}\{A\}$. By the arbitrariness of $B$, we conclude that $\phi_{L}\circ\mathcal{H}\theta_{L}(A)\subseteq\theta_{\mathcal{H}(L)}(A)$.
@@
i.e., to verify $\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)=\mathcal{Q}\eta_{L}(K)$ for any $K\in\mathcal{Q}(L)$, where $\mathcal{Q}\eta_{L}(K)={\uparrow}\{{\downarrow}x:x\in K\}$ and
@@
Let $A$ be an arbitrary element in $\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)$. Then $A\cap K\neq\emptyset$, and thus, there is an $x\in A\cap K$. By the closedness of $A$, we have ${\downarrow}x\subseteq A$, i.e., $A\in{\uparrow}\{{\downarrow}x\}$. So $A\in\mathcal{Q}\eta_{L}(K)$. The arbitrariness of $A$ guarantees that $\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)\subseteq\mathcal{Q}\eta_{L}(K)$.
Now focus on each $x\in K$, we have ${\downarrow}x\cap K\neq\emptyset$. So for any $M\in{\downarrow}\{K\}$, i.e., $K\subseteq M$, ${\downarrow}x\cap M$ always holds. This means $\{{\downarrow}x:x\in K\}\subseteq\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)$; thus, $\mathcal{Q}\eta_{L}(K)\subseteq\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)$.
$\bf{Claim~4:}$ The third diagram in Definition 5.1 commutes, i.e., the following one commutes:
$\mathcal{HQQ}(L)$$\mathcal{QHQ}(L)$$\mathcal{QQH}(L)$$\phi_{\mathcal{Q}(L)}$$\mathcal{Q}\phi_{L}$$\mathcal{HQ}(L)$$\mathcal{QH}(L)$$\mathcal{H}\tau_{L}$$\tau_{\mathcal{H}(L)}$$\phi_{L}$
@@
$\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})=\bigcup{\uparrow}\{\phi_{L}(\mathcal{A}):\forall\mathcal{K}\in\mathbf{A},\mathcal{A}\cap\mathcal{K}\neq\emptyset\}$,
@@
$\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})=\{A\in\mathcal{H}(L):\forall K\in cl(\{\bigcup\mathcal{K}:\mathcal{K}\in\mathbf{A}\}),A\cap K\neq\emptyset\}$.
@@
Consider $\phi_{L}(\mathcal{A})=\{B\in\mathcal{H}(L):\forall M\in\mathcal{A},B\cap M\neq\emptyset\}$, where $\mathcal{A}$ meets each element of $\mathbf{A}$. Suppose that there exists a $B\in\mathcal{H}(L)$ that belongs to $\phi_{L}(\mathcal{A})$ but not to $\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})$, i.e., there is a $K\in cl(\{\bigcup\mathcal{K}:\mathcal{K}\in\mathbf{A}\})$ satisfying that $B\cap K=\emptyset$. It further implies that $K\in\Box(L\setminus B)$. Applying the well-filteredness of $L$ again, we obtain that $\Box(L\setminus B)$ is a Scott open subset of $\mathcal{Q}(L)$; thus, $\Box(L\setminus B)\cap\{\bigcup\mathcal{K}:\mathcal{K}\in\mathbf{A}\}\neq\emptyset$. So we can find a $\mathcal{K}\in\mathbf{A}$ satisfying $\bigcup\mathcal{K}\subseteq L\setminus B$. From which we say that for any $C\in\mathcal{K}$, $C\cap B=\emptyset$. Combining with the fact that $B\in\phi_{L}(\mathcal{A})$, one can conclude that every $C\in\mathcal{K}$ is not in $\mathcal{A}$, that is to say, $\mathcal{K}\cap\mathcal{A}=\emptyset$. However, $\mathcal{A}$ meets every element of $\mathbf{A}$, a contradiction. Thus, each $\phi_{L}(\mathcal{A})$ is contained in $\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})$; accordingly, $\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})\subseteq\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})$.
@@
Take an arbitrary element $A\in\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})$, then for each $\mathcal{K}\in\mathbf{A},A\cap\bigcup\mathcal{K}\neq\emptyset$. So there is an $a_{\mathcal{K}}\in A\cap\bigcup\mathcal{K}$, which further indicates that ${\uparrow}a_{\mathcal{K}}\subseteq K$, i.e., $K\leq{\uparrow}a_{\mathcal{K}}$ for some $K\in\mathcal{K}$. Since $\mathcal{K}$ is an upper set, ${\uparrow}a_{\mathcal{K}}\in\mathcal{K},\forall\mathcal{K}\in\mathbf{A}$. We set $C=cl\{a_{\mathcal{K}}:\mathcal{K}\in\mathbf{A}\}$. Then under the mapping $\mathcal{H}\theta_{L}$, we can obtain a Scott closed subset $cl(\{{\uparrow}c:c\in C\})$ of $\mathcal{Q}(L)$. Next, we verify that $A\in\phi_{L}(cl(\{{\uparrow}c:c\in C\}))$ and $\phi_{L}(cl(\{{\uparrow}c:c\in C\}))\in\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$. First, if there is an $M\in cl(\{{\uparrow}c:c\in C\})$ such that $A\cap M=\emptyset$, then $M\in\Box(L\setminus A)$. Since the well-filteredness of $L$ implies the Scott openness of $\Box(L\setminus A)$, $\Box(L\setminus A)\cap\{{\uparrow}c:c\in C\}\neq\emptyset$. That is, there exists a $c\in C$ such that ${\uparrow}c\subseteq L\setminus A$, then, $c\in L\setminus A$. As $L\setminus A\in\sigma(L)$ and $c\in cl\{a_{\mathcal{K}}:\mathcal{K}\in\mathbf{A}\}$, we can find an $a_{\mathcal{K}}$ such that $a_{\mathcal{K}}\in L\setminus A$. i.e., $a_{\mathcal{K}}\notin A$, obviously, this leads to a contradiction. Thus, $A\in\phi_{L}(cl(\{{\uparrow}c:c\in C\}))$. Then, since for each $a_{\mathcal{K}}\in C$, ${\uparrow}a_{\mathcal{K}}\in cl(\{{\uparrow}c:c\in C\})$ and ${\uparrow}a_{\mathcal{K}}\in\mathcal{K}$, $cl(\{{\uparrow}c:c\in C\})\cap\mathcal{K}\neq\emptyset,\forall\mathcal{K}\in\mathbf{A}$. So $\phi_{L}(cl(\{{\uparrow}c:c\in C\}))\in\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$. As a conclusion, we find an element $\phi_{L}(cl(\{{\uparrow}c:c\in C\}))$ in $\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$ such that $A$ belongs to it, by the arbitrariness of $A$, we get that $\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})\subseteq\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$ holds.
$\bf{Claim~5:}$ The forth diagram in Definition 5.1 commutes, i.e., the following one commutes:

### v5

@@ 5 The double power well-filtered dcpo
In the study of monads, a question always worth considering is the following: given two monads with underlying functors $\mathcal{S}$ and $\mathcal{T}$ on a category $\mathbf{C}$, does $\mathcal{T}\circ\mathcal{S}$ (or $\mathcal{S}\circ\mathcal{T}$) also carry a monad structure? In 1969, J. Beck introduced the concept of distributive laws and showed that their existence provides a sufficient condition for such a composition to form a monad [2].
@@
([2]) Let $\mathcal{S}=(\mathcal{S},\eta^{\mathcal{S}},\mu^{\mathcal{S}})$ and $\mathcal{T}=(\mathcal{T},\eta^{\mathcal{T}},\mu^{\mathcal{T}})$ be two monads on a category $\mathbf{C}$. A distributive law of $\mathcal{S}$ over $\mathcal{T}$ is a natural transformation $l:\mathcal{TS}\rightarrow\mathcal{ST}$ such that the diagrams in Figure 1 commute.
@@
([1]) Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ and $\mathcal{T}^{\prime}=(\mathcal{T}^{\prime},\eta^{\prime},\mu^{\prime})$ be two monads on a category $\mathbf{C}$. A morphism of monads is a natural transformation $\lambda:\mathcal{T}\rightarrow\mathcal{T}^{\prime}$ such that the diagrams in Figure 2 commute.
@@
Particularly, the morphisms of monads are called triple maps by J. Beck in [2]. Meanwhile, J. Beck also proved the following theorem.
@@
([2]) Let $\mathcal{S}=(\mathcal{S},\eta^{\mathcal{S}},\mu^{\mathcal{S}})$, $\mathcal{T}=(\mathcal{T},\eta^{\mathcal{T}},\mu^{\mathcal{T}})$ be two monads on a category $\mathbf{C}$. Then the following statements are equivalent:
@@
      $(\mathcal{TS},\eta_{\mathcal{S}}^{\mathcal{T}}\eta^{\mathcal{S}},\mu)$ is a monad;
@@
    • (c)
      The middle unitary law
      $\mathcal{TS}$$\mathcal{TSTS}$$\mathcal{T}\eta^{\mathcal{S}}_{\mathcal{TS}}\eta^{\mathcal{T}}_{\mathcal{S}}$$\mathcal{TS}$= $\mu$
      i.e., $\mu\circ\mathcal{T}\eta^{\mathcal{S}}_{\mathcal{TS}}\circ\eta^{\mathcal{T}}_{\mathcal{S}}=id_{\mathcal{TS}}$ holds.
From the proposition presented in Section 2 of J. Beck [2], a more general result concerning the morphisms of monads was established by T. Fritz et al.
@@
([7]) Let $\mathcal{S}=(\mathcal{S},\eta_{\mathcal{S}},\mu_{\mathcal{S}})$ and $\mathcal{T}=(\mathcal{T},\eta_{\mathcal{T}},\mu_{\mathcal{T}})$ be two monads on a category $\mathbf{C}$ and $\lambda:\mathcal{S}\rightarrow\mathcal{T}$ a morphism of monads. Then every $\mathcal{T}$-algebra $(A,\alpha)$ can be equipped with an $\mathcal{S}$-algebra structure via $(A,\alpha)\mapsto(A,\alpha\circ\lambda)$. Moreover, a $\mathcal{T}$-algebra homomorphism $f:(A,\alpha)\rightarrow(B,\beta)$ induces an $\mathcal{S}$-algebra homomorphism $f:(A,\alpha\circ\lambda)\rightarrow(B,\beta\circ\lambda)$ in a functorial way.
@@
Let $L$ and $M$ be well-filtered dcpos and $f:L\rightarrow M$ a Scott-continuous map. We need to show that the following diagram commutes:
@@
First, suppose that there exists a $cl(f(A))\in\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})$ but $cl(f(A))\notin\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$. Then we can find an $N_{0}\in cl\{{\uparrow}f(K):K\in\mathcal{A}\}$ satisfying $cl(f(A))\cap N_{0}=\emptyset$. It follows that $N_{0}\in\Box(M\setminus cl(f(A)))$. Since the well-filteredness of $M$ implies that $\Box(M\setminus cl(f(A)))\in\sigma(\mathcal{Q}(M))$, $\Box(M\setminus cl(f(A)))\bigcap\{{\uparrow}f(K):K\in\mathcal{A}\}\neq\emptyset$. So there is a $K_{0}\in\mathcal{A}$ such that ${\uparrow}f(K_{0})\subseteq M\setminus cl(f(A))$, that is to say, ${\uparrow}f(K_{0})\cap cl(f(A))=\emptyset$. Then $K_{0}\cap A=\emptyset$, which obviously contradicts the fact that $A$ meets every element of $\mathcal{A}$. Thus, $\{cl(f(A)):\forall K\in\mathcal{A},A\cap K\neq\emptyset\}\subseteq\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$; therefore, $\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})\subseteq\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$.
Now we prove the reverse inclusion. Let $B\in\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$. Then for each $K\in\mathcal{A}$, $B\cap{\uparrow}f(K)\neq\emptyset$. Because $B$ is a down set, this implies $B\cap f(K)\neq\emptyset$ for all $K\in\mathcal{A}$. We pick $x_{K}\in K$ such that $f(x_{K})\in B\cap f(K)$ and set $A_{0}=cl(\{x_{K}:K\in\mathcal{A}\})$. Then $cl(f(A_{0}))=cl(f(\{x_{K}:K\in\mathcal{A}\}))=cl(\{f(x_{K}):K\in\mathcal{A}\})\subseteq B$ using that $B$ is closed. Meanwhile, $A_{0}\cap K\neq\emptyset$ for each $K\in\mathcal{A}$. So $B\in\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})$; thus, $\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})\subseteq\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})$ holds.
@@
First, suppose for contradiction that there exists a $K\in cl(\{{\uparrow}x:x\in A\})$ such that $A\cap K=\emptyset$. Then $K\subseteq L\setminus A$, i.e., $K\in\Box(L\setminus A)$. By the well-filteredness of $L$, we know $\Box(L\setminus A)\in\sigma(\mathcal{Q}(L))$. So $\Box(L\setminus A)\bigcap\{{\uparrow}x:x\in A\}\neq\emptyset$, that is to say, we can find an $x\in A$ such that ${\uparrow}x\subseteq L\setminus A$, which implies $x\notin A$, a contradiction. So for every $K\in cl(\{{\uparrow}x:x\in A\}),A\cap K\neq\emptyset$. From the definition of $\phi_{L}$, we have $A\in\phi_{L}\circ\mathcal{H}\theta_{L}(A)$; hence, $\theta_{\mathcal{H}(L)}(A)={\uparrow}\{A\}\subseteq\phi_{L}\circ\mathcal{H}\theta_{L}(A)$.
Now we consider the reverse inclusion. Take any $B\in\phi_{L}\circ\mathcal{H}\theta_{L}(A)$. Then for every $x\in A$, $B\cap{\uparrow}x\neq\emptyset$. It means that each $x\in A$ belongs to $B$. Thus $A\subseteq B$, equivalently, $B\in{\uparrow}\{A\}$. By the arbitrariness of $B$, we conclude that $\phi_{L}\circ\mathcal{H}\theta_{L}(A)\subseteq\theta_{\mathcal{H}(L)}(A)$.
@@
i.e., to verify that $\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)=\mathcal{Q}\eta_{L}(K)$ for any $K\in\mathcal{Q}(L)$, where $\mathcal{Q}\eta_{L}(K)={\uparrow}\{{\downarrow}x:x\in K\}$ and
@@
Let $A$ be an arbitrary element in $\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)$; Then $A\cap K\neq\emptyset$; thus, there is an $x\in A\cap K$. By the closedness of $A$, we have ${\downarrow}x\subseteq A$, i.e., $A\in{\uparrow}\{{\downarrow}x\}$. So $A\in\mathcal{Q}\eta_{L}(K)$. The arbitrariness of $A$ guarantees that $\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)\subseteq\mathcal{Q}\eta_{L}(K)$.
Now we focus on the reverse inclusion. For each $x\in K$, we have ${\downarrow}x\cap K\neq\emptyset$. So for any $M\in{\downarrow}\{K\}$, i.e., $K\subseteq M$, ${\downarrow}x\cap M{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\neq\emptyset}$ always holds. This means $\{{\downarrow}x:x\in K\}\subseteq\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)$; thus, $\mathcal{Q}\eta_{L}(K)\subseteq\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)$.
$\bf{Claim~4:}$ The third diagram in Definition 5.1 commutes, i.e., the following one commutes for a well-filtered dcpo $L$:
[…]
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\iota}}_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})=\bigcup{\uparrow}\{\phi_{L}(\mathcal{A}):\forall\mathcal{K}\in\mathbf{A},\mathcal{A}\bigcap\mathcal{K}\neq\emptyset\}$,
@@
$\phi_{L}\circ\mathcal{H}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\iota}_{L}(\mathbf{A})=\{A\in\mathcal{H}(L):\forall K\in cl(\{\bigcup\mathcal{K}:\mathcal{K}\in\mathbf{A}\}),A\cap K\neq\emptyset\}$.
@@
Consider each $\phi_{L}(\mathcal{A})=\{B\in\mathcal{H}(L):\forall M\in\mathcal{A},B\cap M\neq\emptyset\}$, in which $\mathcal{A}\in\mathcal{HQ}(L)$ and meets every element of $\mathbf{A}$. Suppose that there exists a $B\in\mathcal{H}(L)$ belonging to $\phi_{L}(\mathcal{A})$ but not to $\phi_{L}\circ\mathcal{H}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\iota}_{L}(\mathbf{A})$, i.e., there is a $K\in cl(\{\bigcup\mathcal{K}:\mathcal{K}\in\mathbf{A}\})$ satisfying $B\cap K=\emptyset$. Then $K\in\Box(L\setminus B)$. Applying the well-filteredness of $L$, we obtain that $\Box(L\setminus B)$ is a Scott open subset of $\mathcal{Q}(L)$; further, we have $\Box(L\setminus B)\bigcap\{\bigcup\mathcal{K}:\mathcal{K}\in\mathbf{A}\}\neq\emptyset$. So we can find a $\mathcal{K}\in\mathbf{A}$ satisfying $\bigcup\mathcal{K}\subseteq L\setminus B$, which indicates that for any $C\in\mathcal{K}$, $C\cap B=\emptyset$. Combining with the fact that $B\in\phi_{L}(\mathcal{A})$, one can conclude that there is no element of $\mathcal{K}$ belonging to $\mathcal{A}$, that is to say, $\mathcal{K}\bigcap\mathcal{A}=\emptyset$. This contradicts the assumption that $\mathcal{A}$ meets every element of $\mathbf{A}$. Hence, each $\phi_{L}(\mathcal{A})$ is contained in $\phi_{L}\circ\mathcal{H}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\iota}_{L}(\mathbf{A})$; accordingly, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\iota}_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})\subseteq\phi_{L}\circ\mathcal{H}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\iota}_{L}(\mathbf{A})$.
@@
Take an arbitrary element $A\in\phi_{L}\circ\mathcal{H}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\iota}_{L}(\mathbf{A})$. Then for each $\mathcal{K}\in\mathbf{A},A\cap\bigcup\mathcal{K}\neq\emptyset$; hence, we can choose an element $a_{\mathcal{K}}\in A\cap\bigcup\mathcal{K}$. This implies that there exists some $K\in\mathcal{K}$ such that ${\uparrow}a_{\mathcal{K}}\subseteq K$, i.e., $K\leq{\uparrow}a_{\mathcal{K}}$. Since $\mathcal{K}$ is an upper set, ${\uparrow}a_{\mathcal{K}}\in\mathcal{K},\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\forall\mathcal{K}\in\mathbf{A}}$. Set $C=cl\{a_{\mathcal{K}}:\mathcal{K}\in\mathbf{A}\}$. Then the mapping $\mathcal{H}\theta_{L}$ sends the set $C$ to a Scott closed subset $cl(\{{\uparrow}c:c\in C\})$ of $\mathcal{Q}(L)$. We now verify that $A\in\phi_{L}(cl(\{{\uparrow}c:c\in C\}))$ and $\phi_{L}(cl(\{{\uparrow}c:c\in C\}))\in{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})}$. First, suppose that there exists an $M\in cl(\{{\uparrow}c:c\in C\})$ satisfying $A\cap M=\emptyset$. Then $M\in\Box(L\setminus A)$. Since $L$ is well-filtered, $\Box(L\setminus A)$ is a Scott open subset of $\mathcal{Q}(L)$, which implies $\Box(L\setminus A)\bigcap\{{\uparrow}c:c\in C\}\neq\emptyset$. In other words, there exists a $c\in C$ such that ${\uparrow}c\subseteq L\setminus A$; hence, $c\in L\setminus A$. Note that $L\setminus A\in\sigma(L)$ and $c\in cl\{a_{\mathcal{K}}:\mathcal{K}\in\mathbf{A}\}$, so we have $(L\setminus A)\cap\{a_{\mathcal{K}}:\mathcal{K}\in\mathbf{A}\}\neq\emptyset$. Thus we can find an $a_{\mathcal{K}}$ that belongs to $L\setminus A$. i.e., $a_{\mathcal{K}}\notin A$, which obviously leads to a contradiction. Therefore, $A\in\phi_{L}(cl(\{{\uparrow}c:c\in C\}))$. Next, for each $a_{\mathcal{K}}\in C$, we have ${\uparrow}a_{\mathcal{K}}\in cl(\{{\uparrow}c:c\in C\})$ and ${\uparrow}a_{\mathcal{K}}\in\mathcal{K}$ for every $\mathcal{K}\in\mathbf{A}$. It follows that $cl(\{{\uparrow}c:c\in C\})\bigcap\mathcal{K}\neq\emptyset,\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\forall\mathcal{K}\in\mathbf{A}}$. This immediately yields $\phi_{L}(cl(\{{\uparrow}c:c\in C\}))\in{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})}$. In summary, we find an element $\phi_{L}(cl(\{{\uparrow}c:c\in C\}))$ in $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$} such that $A$ belongs to it. By the arbitrariness of $A$, we conclude that $\phi_{L}\circ\mathcal{H}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\iota}_{L}(\mathbf{A})\subseteq{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\iota}_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$.
$\bf{Claim~5:}$ The fourth diagram in Definition 5.1 commutes, i.e., the following one commutes: